OpenStudy (anonymous):

The grades on the last history exam had a mean of 80%. Assume the population of grades on history exams is known to be normally distributed with a standard deviation of 5%. What percent of students earn a score between 70% and 80%? 0.022 0.477 0.50 0.522

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

80 - 70 = 10. The standard deviation is 5. Therefore the grade of 70% is 2 standard deviations below the mean. The empirical rule states that for all normal distributions approximately 95% of the data points lie within the range of plus and minus 2 standard deviations of the mean. In this case only half of the data points in this range are needed, so the required percentage is half of 95%.

OpenStudy (anonymous):

so i would divide that by 1/2

OpenStudy (goformit100):

Sir/Ma'am, A Warm Welcome To Open Study. Sir/Ma'am please show your Working for the question you have posted.

OpenStudy (anonymous):

A survey of 1,000 men and women asked, "Do you earn over $50,000 per year?" The table below shows the responses for males and females: Male Female Total Income over $50,000 485 385 870 Income below $50,000 65 65 130 Total 550 450 1,000 Based on these data, are "being male" and "earning over $50,000" independent events? Yes, P(being male | the person earns over $50,000) ≠ P(being male) Yes, P(being male | the person earns over $50,000) = P(being male) No, P(being male | the person earns over $50,000) ≠ P(being male) No, P(being male | the person earns over $50,000) = P(being male) @ganeshie8

OpenStudy (anonymous):

Independent events: P(A and B) = P(A)P(B) P(male, >50k) = 485/1000 = .485 P(male) = 550/1000 = .550 P(>50k) = 870/1000 = .870 P(male)P(>50k) = .4785 not 0.485 thus not independent The options seem to use another criterion, however.

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

I guess if independent then P(being male and >50k/yr) = P(being male) because making >5ok would not make you more or less likely to be a male.

OpenStudy (anonymous):

u should give me a medal ;P

OpenStudy (anonymous):

i think its C